# Importance sampling for Sobol’ indices estimation

### v2

@@ 1 Introduction
To quantify this impact, a large variety of sensitivity indices have been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [34, 36], others are based on discrepancy measures between probability distributions [4, 5, 10], while others use contrast-based functionals or Cramér–von Mises distance [20, 24]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [42], they are particularly well suited for scalar outputs, and can be generalised to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [27], Sobol’ indices measure the contribution of each group of inputs to the output variance. The first-order Sobol’ indices quantify the main effects of each input, while the total-order Sobol’ indices also account for both main and interaction effects. Hence, they are variance-based measures that capture the contribution of inputs to the overall variability of the model output.
@@
With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, as it requires two evaluations of the model for each realization of $X$, which is often prohibitive. When the context allows it, using this particular design can considerably simplify estimation, but the number of code calls often makes it prohibitive. By contrast, given-data methods are increasingly favored because they operate from a provided input–output sample only. This naturally raises the question of efficiency limits in the given-data setting.
Among them, several estimators achieve the asymptotic efficiency lower bound. The expression of the efficiency bound for $\eta$ was first derived in [18] for a continuous input $X$, from the efficient influence function associated with the parameter. The same work also proposed an efficient estimator of a truncated version of $\eta$. Since then, several methods have been shown to be asymptotically efficient in the given-data setting. For first-order indices, the kernel estimator introduced in [11] achieves the asymptotic optimal variance bound for estimating $\eta$ (see Theorems 3.4 and 3.5). This result was initially established for first-order Sobol’ indices and later extended to higher-order ones under suitable regularity conditions [12] (see Theorems 4.1 and 4.2). These estimators, however, remain difficult to implement in practice. In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalized to attain the efficiency bound in [31] (Proposition 3.3). The question of efficiency bounds has been thoroughly discussed in [30]. In this work, we mainly focus on these efficient estimators.
@@
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is the most challenging term to estimate accurately. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution, which makes it possible to characterize optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
@@ 2 Main notation
This estimator achieves zero variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere. This ideal distribution cannot be used in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as nonparametric methods [49] or parametric approaches like the cross-entropy method [38, 14]. The efficiency of IS [37] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values.
The next section extends the classical IS framework to characterize the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
@@ 3.1 Input distribution change via importance sampling
We introduce the importance sampling framework by allowing a modification of the input distribution of the random vector $\mathbf{X}\in\mathbb{R}^{k}$, while keeping the noise variable $\mathbf{W}$ fixed. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$. For any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\mathbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x}_{u})$ its marginal density. We introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
@@ 3.2 Optimal asymptotic variance expression using importance sampling
$\boxed{\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.}$ (11)
@@ 3.3 Optimal distribution: sequential construction
To characterize the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations. Following the classical Lagrangian minimisation framework, one can formally derive the stationary condition satisfied by $q^{*}$ under the normalisation constraint. In general, verifying that this stationary point corresponds to a minimum would require checking second-order optimality conditions. However, as mentioned by Owen [34, End Notes, p. 38], the Cauchy–Schwarz inequality is sufficient to establish global minimality in this setting. For simplicity, we rely on the following key variational lemma, a standard result in importance sampling theory.
@@
3.4 Zero variance optimal distribution
@@
In general, this distribution does not yield a zero variance estimator, since a direct plug-in calculation shows
@@
Remark. When the model $f(\mathbf{x})$ may take both positive and negative values, the theoretical zero variance density derived above is no longer a valid probability distribution. A practical alternative consists in separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution. This idea, sometimes referred to as positivisation in the importance sampling literature [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
@@
Two auxiliary densities $q_{+}$ and $q_{-}$ are then defined to be positive wherever $pf_{+}$ and $pf_{-}$ are positive, respectively. Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, preserving unbiasedness while reducing variance in practice.
@@
$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
In theory, the density $q^{*}(\mathbf{x})$ yields a zero variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it. Adaptive procedures can be employed, including nonparametric methods [49] and parametric ones such as the cross-entropy method [38, 14].
@@ 5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution that is different from the reference one. In particular, we show that by choosing suitable auxiliary densities, the asymptotic variance of efficient estimators can be further reduced, as changing the sampling distribution modifies the efficiency bound itself. In some cases, zero variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. In realistic settings where model evaluations are available, one may nevertheless approximate these optimal densities, which naturally leads to significant variance reduction. We also show that combining our method with the rank-based estimator [22] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [32]. The first limitation of our present work is that the optimal densities depend on the code, so additional computational effort is required. Another limitation is that we have proved the existence of one optimal distribution for every Sobol’ index, so it may be computationally expensive to estimate all these optimal densities, which increases the numerical effort. In a related direction, [15] investigate how several expectations under different target distributions can be estimated from a single Monte Carlo sample, using an adaptive importance sampling scheme with shared control variates. This suggests that multi–target designs could also be explored to construct sampling densities efficient for several Sobol’ quantities simultaneously. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
@@ Acknowledgement
Derivation of the zero variance formula We start from the general expression:
@@ Appendix B. Optimal sampling densities for the rank estimator
$q^{*}_{\mathrm{rank}}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$

### v3

# arXiv:2507.05958v3[math.ST] 01 Mar 2026

@@ 1 Introduction
To quantify this impact, a large variety of sensitivity indices have been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [34, 37], others are based on discrepancy measures between probability distributions [4, 5, 10], while others use contrast-based functionals or Cramér–von Mises distance [20, 24]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [43], they are particularly well suited for scalar outputs, and can be generalised to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [27], Sobol’ indices measure the contribution of each group of inputs to the output variance. The first-order Sobol’ indices quantify the main effect of each input, while the total-order Sobol’ indices also account for interaction effects. Hence, they are variance-based measures that capture the contribution of inputs to the overall variability of the model output.
@@
With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, since the number of model evaluations required to maintain a given accuracy for all indices increases linearly with the number $k$ of input variables. When the context allows it, using this particular design can considerably simplify estimation, but the number of code calls often makes it prohibitive. By contrast, given-data methods are increasingly favored because they operate from a provided input–output sample only. This naturally raises the question of efficiency limits in the given-data setting.
Among them, several estimators achieve the asymptotic efficiency lower bound. The expression of the efficiency bound for $\eta$ was first derived in [18] for a continuous input $X$, from the efficient influence function associated with the parameter. The same work also proposed an efficient estimator of a truncated version of $\eta$. Since then, several methods have been shown to be asymptotically efficient in the given-data setting. For first-order indices, the kernel estimator introduced in [11] achieves the asymptotic optimal variance bound for estimating $\eta$ (see Theorems 3.4 and 3.5). This result was initially established for first-order Sobol’ indices and later extended to higher-order ones under suitable regularity conditions [12] (see Theorems 4.1 and 4.2). These estimators, however, remain difficult to implement in practice. In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalised to attain the efficiency bound in [31] (Proposition 3.3). The question of efficiency bounds has been thoroughly discussed in [30]. In this work, we mainly focus on these efficient estimators.
@@
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators is the most challenging term to estimate accurately. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution,which makes it possible to characterise optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
@@ 2 Main notation
This estimator achieves zero-variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere. This ideal distribution cannot be used in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as nonparametric methods [50] or parametric approaches like the cross-entropy method [39, 14]. The efficiency of IS [38] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values.
The next section extends the classical IS framework to characterise the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
@@ 3.1 Input distribution change via importance sampling
We introduce the importance sampling framework by allowing a modification of the input distribution of the random vector $\mathbf{X}\in\mathbb{R}^{k}$, while keeping the noise variable $\mathbf{W}$ fixed. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$. For any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\mathbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x}_{u})$ its marginal density. We introduce an auxiliary joint density $q(\mathbf{x})$, factorised as
@@ 3.2 Optimal asymptotic variance expression using importance sampling
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{q}\Big[4\,\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{2}\,\mathrm{Var}_{q}(Z_{u}\mid\mathbf{X}_{u})+\big(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\big)^{4}\Big]-\eta_{u}^{2}.$ (11)
@@ 3.3 Optimal distribution: sequential construction
To characterise the optimal distribution $q^{*}$, a principled approach consists in using the calculus of variations. Following the classical Lagrangian minimisation framework, one can formally derive the stationary condition satisfied by $q^{*}$ under the normalisation constraint. In general, verifying that this stationary point corresponds to a minimum would require checking second-order optimality conditions. However, as mentioned by Owen [35, Chapter 9, End Notes (“Calculus of variations”), p. 38], the Cauchy–Schwarz inequality is sufficient to establish global minimality in this setting. For simplicity, we rely on the following key variational lemma, a standard result in importance sampling theory.
@@
3.4 zero-variance optimal distribution
@@
For instance, this distribution does not yield a zero-variance estimator, since a direct plug-in calculation shows
@@
Remark. When the model $f(\mathbf{x})$ takes both positive and negative values, the associated optimal variance is strictly positive, so the zero-variance property cannot be achieved. If one nevertheless wishes to construct a zero-variance estimator, a classical alternative consists of separating the positive and negative contributions of $f$, so that each part can be sampled using its own importance distribution. This idea, referred to as positivisation [34], relies on defining two auxiliary densities associated with the positive and negative parts of $f$.
@@
Two auxiliary densities $q_{+}$ and $q_{-}$ are then defined to be positive wherever $pf_{+}$ and $pf_{-}$ are positive, respectively. Independent samples $\mathbf{X}_{i,+}\sim q_{+}$ and $\mathbf{X}_{i,-}\sim q_{-}$ are then combined with opposite signs, yielding an unbiased estimator. In the ideal case where each component is sampled according to its fully adapted optimal density, the resulting estimator achieves zero variance.
@@
$q_{u}^{*}(\mathbf{x})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (18)
In theory, the density $q^{*}(\mathbf{x})$ yields a zero-variance estimator, but this ideal case cannot be achieved in practice since the optimal distribution depends on the unknown target quantity $\eta_{u}$, which brings us back to the classical importance sampling problem of approximating it. Adaptive procedures can be employed, including nonparametric methods [50] and parametric ones such as the cross-entropy method [39, 14].
@@ 5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution that is different from the reference one. In particular, we show that by choosing suitable auxiliary densities, the asymptotic variance of efficient estimators can be further reduced, as changing the sampling distribution modifies the efficiency bound itself. In some cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. In realistic settings where model evaluations are available, one may nevertheless approximate these optimal densities, which naturally leads to significant variance reduction. We also show that combining our method with the rank-based estimator [22] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [32]. The first limitation of our present work is that the optimal densities depend on the code, so additional computational effort is required. Another limitation is that we have proved the existence of one optimal distribution for every Sobol’ index, so it may be computationally expensive to estimate all these optimal densities, which increases the numerical effort. In a related direction, [15] investigate how several expectations under different target distributions can be estimated from a single Monte Carlo sample, using an adaptive importance sampling scheme with shared control variates. This suggests that multi–target designs could also be explored to construct sampling densities efficient for several Sobol’ quantities simultaneously. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
@@ Acknowledgement
Derivation of the zero-variance formula We start from the general expression:
@@ Appendix B. Optimal sampling densities for the rank estimator
$q_{u}^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\,q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$
@@
Deterministic case ($Y=f(\mathbf{X})$).
Assume that $f\geq 0$ and that there is no stochastic noise, i.e. $Y=f(\mathbf{X})$. Then $\phi^{2}(\mathbf{x})=\mathbb{E}[Y^{2}\mid\mathbf{X}=\mathbf{x}]=f^{2}(\mathbf{x})$ and thus $\phi(\mathbf{x})=f(\mathbf{x})$. In particular,
$m(\mathbf{x}_{u})=\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}],\qquad\nu_{p}(\mathbf{x}_{u})=\operatorname{Var}_{p}(f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}).$
The optimal conditional density becomes
$q^{*}_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]}=\frac{p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})\,f(\mathbf{x})}{m(\mathbf{x}_{u})}.$
Moreover, under $q^{*}_{-u\mid u}$ we have
$\nu_{q^{*}}(\mathbf{x}_{u})=\big(\mathbb{E}_{p}[f(\mathbf{X})\mid\mathbf{X}_{u}=\mathbf{x}_{u}]\big)^{2}-m^{2}(\mathbf{x}_{u})=0,$
so that $S_{B}(\mathbf{x}_{u})=m^{4}(\mathbf{x}_{u})$ and hence
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m^{2}(\mathbf{x}_{u})}{\eta_{u}}.$
Therefore, the resulting optimal sampling density is
$q_{u}^{*}(\mathbf{x})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m(\mathbf{x}_{u})}{\eta_{u}}.$